OpenStudy (anonymous):

find the solution of this system of equations -6x+3y=-39 -6x-3y=-21

jhonyy9 (jhonyy9):

-6x +3y = -39 -6x -3y = -21 ------------------ -12x 0 = - 60 x = - 60 /(-12) = 5 x =5 -6*5 -3y = -21 -30 -3y = -21 -3y = 9 y = 9/(-3) y = -3 x = 5 y = - 3

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

was my pleasure